# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ Appendix D Proof of Lemma 10
The state $\rho_{\bm{i}}$ is equivalent to $\bigotimes_{l\in[k]}\rho_{l}^{\otimes n_{l}(\bm{i})}$ up to permutation. Note that the permutation, which is composed of the SWAP operations among subsystems, is a $G$-covariant operation. For a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$, $n_{l}(\bm{l})\leq N(p_{l}+\delta)$ holds for all $l\in[k]$. Therefore, a state $\rho_{\bm{i}}$ for a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$ can be created by first preparing $\bigotimes_{l\in[k]}\rho_{l}^{\otimes\lfloor N(p_{l}+\delta)\rfloor}$ via a $G$-covariant operation, discarding some copies if needed, and then operating a permutation. For each $l\in[k]$, for $r_{l}$ satisfying $r_{l}>(p_{l}+\delta)\mathcal{A}_{\mathrm{c}}(\rho_{l})$, it holds
@@
Appendix F Proof of an upper bound on the asymmetry cost
In this section, we present the proof of Proposition 14, which provides an upper bound on the asymmetry cost. It is worth emphasizing that, unlike the result in the previous section, Proposition 14 does not impose any constraint on the symmetry subgroups in the ensemble decomposition. Intuitively, this relaxation is possible because the symmetry subgroup constraint is only relevant for the estimation step in the estimation-and-conversion strategy introduced in Sec.V, while the conversion step succeeds as long as the monotonicity condition of QGTs is satisfied. Below, we carefully modify the argument of typical sequences presented in the previous section, and use it to extend the proof of Proposition 6, thereby establishing Proposition 14. Following the convention in the main text, we define $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ for a unitary representation $U$ of $G$. Note that, with a slight abuse of notation, we also write $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ for $\bm{\theta}\in\mathbb{R}^{\dim G}$.
@@ Appendix F Proof of an upper bound on the asymmetry cost
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\rho$ be a state on $\mathcal{H}$ and $\rho$ be a state on $\mathcal{H}^{\prime}$. Consider an ensemble $\{p_{i},\sigma_{i}\}_{i\in[k]}$ such that $\sigma=\sum_{i}p_{i}\sigma_{i}$, where $k]\coloneqq\{1,\cdots,k\}$. Suppose that for each $i$, for some finite $r_{i}>0$, there exists a sequence of quantum channels $\mathcal{E}_{N}^{(i,g)}$ such that
$\delta_{N}^{(i)}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(i,g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil r_{i}N\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{i,g}\right)^{\otimes N}\right)$ (413)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}^{(i)}(g,\bm{u})=0$ for $\epsilon^{\prime}\in(0,\epsilon)$, where $\rho_{g}\coloneqq\mathcal{U}_{g}(\rho)$ and $\sigma_{i,g}\coloneqq\mathcal{U}_{g}^{\prime}(\sigma_{i})$. Then there exists a sequence of quantum channels $\mathcal{E}_{N}^{(g)}$ and $\epsilon^{\prime}\in(0,1)$ such that
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)$ (414)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{g})\right)^{\otimes N},\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime\otimes N}\circ\mathcal{U}_{g}^{\prime\otimes N}\right)\left(\tilde{\sigma}_{N}\right)\right)=0.$ (419)
@@
$T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}T\left(\mathcal{E}^{(l,g)}_{N_{l}}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})^{\otimes N_{l}}\right)$
$=\sum_{l\in[k]}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\frac{\sqrt{N_{l}}}{\sqrt{N}}\right).$ (421)
@@
$\|\bm{u}\|<N^{\epsilon}\implies\left\|\bm{u}\frac{\sqrt{N_{l}}}{\sqrt{N}}\right\|<N_{l}^{\epsilon}.$ (422)
@@
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right)$ (423)
@@
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{D}_{N}^{(\bm{i})}\circ\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes n_{l}(\bm{i})}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right).$ (424)
Denoting by $\mathcal{P}_{\bm{i}}$ a channel that swaps subsystems so that $\mathcal{P}^{(\bm{i})}(\bigotimes_{l\in[k]}\sigma_{l}^{\otimes n_{l}(\bm{i})})=\sigma_{\bm{i}}$, which commutes with $\mathcal{U}^{\otimes N}_{\bm{u}}\circ\mathcal{U}^{\otimes N}_{g}$, we get
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right)$ (425)

### v4

@@ D.2.5 Proof of Lemma 27
D.3 Asymptotic conversion error
We study the asymptotic behavior of the conversion error in our channels. While we do not aim to determine the optimal error rate, our results will provide a basis for future higher-order analysis. In this subsection, we repeatedly use the Fuchs–van de Graaf inequalities, i.e.,
$1-\sqrt{\mathrm{Fid}(\rho,\sigma)}\leq T(\rho,\sigma)\leq\sqrt{1-\mathrm{Fid}(\rho,\sigma)}.$ (402)
In what follows, we fix some $\epsilon\in(0,1/9)$. We first analyze the errors in the conversion from $\{\psi^{\otimes N}_{u/\sqrt{N}}\}_{N}$ to $\{\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\}_{N}$. For this purpose, we investigate each of the following steps (i)–(iv):
[…]
Here, we remind the readers of our notations: $\tilde{\psi}_{\theta}=e^{\mathrm{i}\theta^{\mu}\tilde{X}_{\mu}}\psi e^{-\mathrm{i}\theta^{\nu}\tilde{X}_{\nu}}$ and $\tilde{\phi}_{\theta}=e^{\mathrm{i}\theta^{\mu}\tilde{X}^{\prime}_{\mu}}\psi e^{-\mathrm{i}\theta^{\nu}\tilde{X}^{\prime}_{\nu}}$ for $\theta\in\mathbb{R}^{\dim G}$ with $\tilde{X}_{\mu}=(I-\psi)X_{\mu}\psi+\psi X_{\mu}(I-\psi)$ and $\tilde{X}^{\prime}_{\mu}=(I-\phi)X^{\prime}_{\mu}\phi+\phi X_{\mu}^{\prime}(I-\phi)$.
Steps (i) and (iv): From Lemma 30 and Eq. (274) in its proof, we have
$\left|\mathrm{Fid}\left(\psi_{u/\sqrt{N}},\tilde{\psi}_{u/\sqrt{N}}\right)-1\right|$ $\leq a_{1}N^{-3/2+3\epsilon}$ (404)
$\left|\mathrm{Fid}\left(\phi_{u/\sqrt{N}},\tilde{\phi}_{u/\sqrt{N}}\right)-1\right|$ $\leq a_{2}N^{-3/2+3\epsilon}$ (405)
with some finite non-negative constants $a_{1}$ and $a_{2}$ independent of $N$. Since the fidelity is multiplicative, and is less than or equal to one, we obtain
$\mathrm{Fid}\left(\psi_{u/\sqrt{N}}^{\otimes N},\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}\right)$ $=\mathrm{Fid}\left(\psi_{u/\sqrt{N}},\tilde{\psi}_{u/\sqrt{N}}\right)^{N}\geq\left(1-a_{1}N^{-3/2+3\epsilon}\right)^{N}=1-O(N^{-1/2+3\epsilon})$ (406)
$\mathrm{Fid}\left(\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor},\tilde{\phi}_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)$ $=\mathrm{Fid}\left(\phi_{u/\sqrt{N}},\tilde{\phi}_{u/\sqrt{N}}\right)^{\lfloor rN\rfloor}\geq\left(1-a_{2}N^{-3/2+3\epsilon}\right)^{\lfloor rN\rfloor}=1-O(N^{-1/2+3\epsilon})$ (407)
where we used $\epsilon<1/9<1/6$. Therefore, we find
$\sup_{\|u\|<N^{\epsilon}}T\left(\psi_{u/\sqrt{N}}^{\otimes N},\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}\right)$ $\leq\inf_{\|u\|<N^{\epsilon}}\sqrt{1-\mathrm{Fid}\left(\psi_{u/\sqrt{N}}^{\otimes N},\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}\right)}=O(N^{-\frac{1}{4}(1-6\epsilon)}),$ (408)
$\sup_{\|u\|<N^{\epsilon}}T\left(\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor},\tilde{\phi}_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)$ $\leq\inf_{\|u\|<N^{\epsilon}}\sqrt{1-\mathrm{Fid}\left(\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor},\tilde{\phi}_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)}=O(N^{-\frac{1}{4}(1-6\epsilon)}).$ (409)
Step (ii): For $r$ satisfying $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$, in Lemma 30 and its proof, we have constructed a quantum channel $\mathcal{E}$ satisfying Eq. (328), i.e.,
$\left|\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{u/\sqrt{N}}),\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}\right)-1\right|\leq a_{3}N^{-3/2+3\epsilon}$ (410)
for a finite non-negative constant $a_{3}$ independent of $N$. Therefore, as in Steps (i) and (iv), we obtain
$\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}\left(\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}\right),\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\right)$ $\leq\inf_{\|u\|<N^{\epsilon}}\sqrt{1-\mathrm{Fid}\left(\mathcal{E}\left(\tilde{\psi}_{u/\sqrt{N}}\right),\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}\right)^{N}}=O(N^{-\frac{1}{4}(1-6\epsilon)}).$ (411)
Step (iii): The sequences of channels $\{\mathcal{T}_{N}\}_{N}$ and $\{\mathcal{S}_{N}\}_{N}$ in Eqs.(292) and (293) are provided in Theorem 3.3 of [73]. The conversion errors are shown to be polynomial in $N$, i.e.,
[…]
holds for some $\kappa>0$.
Since the term arising from Eq. (289) is of order $O(N^{-1/2})$ and hence subdominant, we get
[…]
Consequently, we get
$\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\left(\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\right),\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)=O\left(N^{-\kappa}\right).$ (415)
From Eqs. (408), (409), (411), and (415), we obtain
$\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\circ\mathcal{E}^{\otimes N}\left(\psi_{u/\sqrt{N}}^{\otimes N}\right),\phi^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)=O\left(N^{-\kappa^{\prime}}\right),\quad\kappa^{\prime}\coloneqq\min\left\{\kappa,\frac{1}{4}(1-6\epsilon)\right\}.$ (416)
We remark that the conversion channels $\mathcal{S}_{\lfloor rN\rfloor}$ and $\mathcal{T}_{N}$ are constructed in the proof of Theorem 3.3 of [73], while the channel $\mathcal{E}$ is constructed in Lemma 30.
From these sequences of channels, the sequence of channels $\{\mathcal{E}_{N}^{(g)}\}$ is constructed in the proof of Lemma 27, which satisfies
$\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}\delta_{N}(g,u)$ $=\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)=O\left(N^{-\kappa^{\prime}}\right),$ (417)
where $\epsilon_{1}$ is an arbitrarily small positive constant. Therefore, for the channels $\{\mathcal{E}_{N}\}$ defined in Eq. (235), from Eqs. (238), (239) and (390), we obtain
$\inf_{g\in G}\mathrm{Fid}\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)\geq p^{\mathrm{succ.}}(N^{-1/2+\epsilon})\times f_{N}=\left(1-O\left(N^{-\kappa^{\prime}}\right)\right)^{2},$ (418)
implying that
$\sup_{g\in G}T\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)=O\left(N^{-\frac{1}{2}\kappa^{\prime}}\right).$ (419)
Following the arguments around Eq. (242), for an arbitrarily small $\delta>0$, by using the channel $\Lambda_{N}$ discarding $\lfloor rN^{\prime}\rfloor-\lfloor(r-\delta)N\rfloor$ copies of the output system, we find
$\sup_{g\in G}T\left(\Lambda_{N}\circ\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor(r-\delta)N\rfloor}\right)=O\left(N^{-\frac{1}{2}\kappa^{\prime}}\right).$ (420)
Finally, we introduce
$\mathcal{E}_{N}^{\prime}\coloneqq\int\mathrm{d}\mu_{G}(g)\,\mathcal{U}_{g^{-1}}^{\prime\otimes\lfloor(r-\delta)N\rfloor}\circ\Lambda_{N}\circ\mathcal{E}_{N}\circ\mathcal{U}_{g}^{\otimes N}.$ (421)
From Lemma 4, $\{\mathcal{E}_{N}^{\prime}\}_{N}$ is a sequence of $G$-covariant channels such that
$T\left(\mathcal{E}_{N}^{\prime}\left(\psi^{\otimes N}\right),\phi^{\otimes\lfloor(r-\delta)N\rfloor}\right)=O\left(N^{-\frac{1}{2}\kappa^{\prime}}\right).$ (422)
Therefore, with this sequence of conversion channels $\{\mathcal{E}_{N}^{\prime}\}_{N}$, the conversion error in trace distance decays polynomially with $N$.
Adopting the result of [73], we see that the coefficient $\kappa^{\prime}$ is determined by $\kappa$, which characterizes the conversion error in QLAN. This implies that an improvement of the QLAN error rate directly translates into an improvement of our conversion error. To demonstrate this, we use Lemma C.6 of [73], which states that
[…]
$=O\left(N^{-\frac{1}{4}+\Delta}+N^{-\frac{1}{2}+\zeta}+N^{-\frac{1}{2}+\zeta+\eta}+N^{\frac{9\eta-2}{24}}+N^{-\frac{1}{4}+\frac{3\epsilon}{2}+\Delta}+N^{-\frac{1}{2}+\frac{\alpha}{2}+\frac{\eta}{2}}+N^{-\frac{1}{2}+\frac{3\eta}{2}}+N^{-\frac{\Delta}{2}}\right)$ (423)
where $\zeta$, $\Delta$, $\alpha$, $\eta$ are constants, which can be arbitrarily chosen as long as they satisfy
$0<\zeta<\frac{1}{4},\quad 0<\Delta,\quad\frac{1}{2}+\zeta<\alpha,\quad\eta<\frac{2}{9},\quad 0<\epsilon<\frac{\eta-\Delta}{2}.$ (424)
The same bound applies $\mathcal{S}_{N}$ as noted in Eq. (C.9) of [73]. For example, let $\epsilon_{0}\in(4\epsilon/7,4/63)$ and choose
$\zeta=\epsilon_{0},\quad\Delta=\frac{2}{21}-\frac{3}{2}\epsilon_{0},\quad\alpha=\frac{1}{2}+2\zeta=\frac{1}{2}+2\epsilon_{0},\quad\eta=\frac{2}{21}+2\epsilon_{0},$ (425)
which satisfy the above constraints. Substituting these parameters yields
[…]
Hence, $\kappa$ can be chosen arbitrarily close to $\frac{1}{21}$.
@@ D.3 Asymptotic conversion error
The state $\rho_{\bm{i}}$ is equivalent to $\bigotimes_{l\in[k]}\rho_{l}^{\otimes n_{l}(\bm{i})}$ up to permutation. Note that the permutation, which is composed of the SWAP operations among subsystems, is a $G$-covariant operation. For a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$, $n_{l}(\bm{i})\leq N(p_{l}+\delta)$ holds for all $l\in[k]$. Therefore, a state $\rho_{\bm{i}}$ for a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$ can be created by first preparing $\bigotimes_{l\in[k]}\rho_{l}^{\otimes\lfloor N(p_{l}+\delta)\rfloor}$ via a $G$-covariant operation, discarding some copies if needed, and then operating a permutation. For each $l\in[k]$, for $r_{l}$ satisfying $r_{l}>(p_{l}+\delta)\mathcal{A}_{\mathrm{c}}(\rho_{l})$, it holds
@@
Appendix F Proof of an upper bound on the asymmetry cost (Proposition 7)
In this section, we present the proof of Proposition 7, which provides an upper bound on the asymmetry cost. It is worth emphasizing that, unlike the result in the previous section, Proposition 7 does not impose any constraint on the symmetry subgroups in the ensemble decomposition. Intuitively, this relaxation is possible because the symmetry subgroup constraint is only relevant for the estimation step in the estimation-and-conversion strategy introduced in Sec. D.1, while the conversion step succeeds as long as the monotonicity condition of QGTs is satisfied. Below, we carefully modify the argument of typical sequences presented in the previous section, and use it to extend the proof of Proposition 24, thereby establishing Proposition 7. Following the convention in the main text, we define $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ for a unitary representation $U$ of $G$. Note that, with a slight abuse of notation, we also write $\mathcal{U}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ for $\theta\in\mathbb{R}^{\dim G}$.
@@ Appendix F Proof of an upper bound on the asymmetry cost (Proposition 7)
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\rho$ be a state on $\mathcal{H}$ and $\sigma$ be a state on $\mathcal{H}^{\prime}$. Consider an ensemble $\{p_{i},\sigma_{i}\}_{i\in[k]}$ such that $\sigma=\sum_{i}p_{i}\sigma_{i}$, where $k]\coloneqq\{1,\cdots,k\}$. Suppose that for each $i$, for some finite $r_{i}>0$, there exists a sequence of quantum channels $\mathcal{E}_{N}^{(i,g)}$ such that
$\delta_{N}^{(i)}(g,u)\coloneqq T\left(\mathcal{E}^{(i,g)}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil r_{i}N\rceil}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}\left(\sigma_{i,g}\right)^{\otimes N}\right)$ (439)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}\delta_{N}^{(i)}(g,u)=0$ for $\epsilon^{\prime}\in(0,\epsilon)$, where $\rho_{g}\coloneqq\mathcal{U}_{g}(\rho)$ and $\sigma_{i,g}\coloneqq\mathcal{U}_{g}^{\prime}(\sigma_{i})$. Then there exists a sequence of quantum channels $\mathcal{E}_{N}^{(g)}$ and $\epsilon^{\prime}\in(0,1)$ such that
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)$ (440)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}T\left(\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\sigma_{g})\right)^{\otimes N},\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\prime\otimes N}\circ\mathcal{U}_{g}^{\prime\otimes N}\right)\left(\tilde{\sigma}_{N}\right)\right)=0.$ (445)
@@
$T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}T\left(\mathcal{E}^{(l,g)}_{N_{l}}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\sigma_{l,g})^{\otimes N_{l}}\right)$
$=\sum_{l\in[k]}\delta_{N_{l}}^{(l)}\left(g,u\frac{\sqrt{N_{l}}}{\sqrt{N}}\right).$ (447)
@@
$\|u\|<N^{\epsilon}\implies\left\|u\frac{\sqrt{N_{l}}}{\sqrt{N}}\right\|<N_{l}^{\epsilon}.$ (448)
@@
$\sup_{g\in G}\sup_{\|u\|<N^{\epsilon^{\prime}}}T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|u\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,u\right)$ (449)
@@
$\sup_{g\in G}\sup_{\|u\|<N^{\epsilon^{\prime}}}T\left(\mathcal{D}_{N}^{(\bm{i})}\circ\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes n_{l}(\bm{i})}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|u\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,u\right).$ (450)
Denoting by $\mathcal{P}_{\bm{i}}$ a channel that swaps subsystems so that $\mathcal{P}^{(\bm{i})}(\bigotimes_{l\in[k]}\sigma_{l}^{\otimes n_{l}(\bm{i})})=\sigma_{\bm{i}}$, which commutes with $\mathcal{U}^{\otimes N}_{u}\circ\mathcal{U}^{\otimes N}_{g}$, we get
$\sup_{g\in G}\sup_{\|u\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{u}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|u\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,u\right)$ (451)